# Integrating anatomy and electrophysiology in the healthy human heart: Insights from biventricular statistical shape analysis using universal coordinates

### v1

@@ Abstract
A cardiac digital twin is a virtual replica of a patient-specific heart, mimicking its anatomy and physiology. A crucial step of building a cardiac digital twin is anatomical twinning, where the computational mesh of the digital twin is tailored to the patient-specific cardiac anatomy. In a number of studies, the effect of anatomical variation on clinically relevant functional measurements like electrocardiograms (ECGs) is investigated, using computational simulations. While such a simulation environment provides researchers with a carefully controlled ground truth, the impact of anatomical differences on functional measurements in real-world patients remains understudied. In this study, we develop a biventricular statistical shape model and use it to quantify the effect of biventricular anatomy on ECG-derived and demographic features, providing novel insights for the development of digital twins of cardiac electrophysiology. To this end, a dataset comprising high-resolution cardiac CT scans from 271 healthy individuals, including athletes, is utilized. Furthermore, a novel, universal, ventricular coordinate-based method is developed to establish lightweight shape correspondence. The performance of the shape model is rigorously established, focusing on its dimensionality reduction capabilities and the training data requirements. Additionally, a comprehensive synthetic cohort is made available, featuring ready-to-use biventricular meshes with fiber structures and anatomical region annotations. These meshes are well-suited for electrophysiological simulations.
@@ 1.1 Cardiac digital twins
Regardless of the specific aspect(s) of the cardiovascular system that a model targets, the anatomy of the individual needs to be taken into account. As such, defining a computational mesh representing this anatomy is a common step in the construction of all types of CDTs. This discrete, multi-label mesh is typically derived from cardiac computed tomography (CT) or magnetic resonance (MR) imaging of the subject. It delineates the computational domains for discretizing and numerically solving the PDEs governing the system. The process of creating a mesh that matches subject-specific anatomy is known as anatomical twinning [11]. Cardiac statistical shape modeling, which captures population-level anatomical variations, closely relates to anatomical twinning, both in methodology and applications, offering a powerful means to study and represent individual differences within a broader population framework [12, 13, 14, 15, 16].
@@ 1.2 Statistical shape modeling
(3) Statistical analysis: A cardiac SSM offers valuable, quantitative insights into heart shape variability, which can in turn be linked to other health-related data, such as electrophysiological measurements and demographic information, through statistical analyses [15, 16]. As such, cardiac SSM are valuable tools for advancing our fundamental understanding of cardiac anatomy, notably in relation to other clinically relevant features, such as electrocardiograms (ECGs).
@@ 1.3 Related work
Previous work in cardiac SSM development provides meaningful insights into cardiac anatomy of various scopes (biventricular anatomy in [13, 14], four-chamber cardiac anatomy in [15], and bi-atrial anatomy in [16]). However, some limitations can be identified. The spatial resolution of the input data is often limited compared to the clinical state-of-the-art [25, 13, 14, 16]. Furthermore, some studies include only a limited and possibly insufficient number of subjects in their shape analysis, such as 19 shapes in [15] and 47 shapes in [16]. Such limitations are often related to the lack of automatic workflows for the generation of training geometries, resulting in time-consuming manual processes, especially in the segmentation step [14, 15, 16]. Moreover, none of the studies leverage the recent advancements in the field of cardiac modeling, specifically designed for registration purposes, i.e. the implementation of universal cardiac coordinate systems [26, 27, 11]. Additionally, although cardiac SSM studies usually include the release of synthetic cohorts, many of these cohorts only feature surface meshes [13, 14], and as such are not suitable for direct use in downstream applications, such as finite element simulations. Notable exceptions are [15] and [16], which include the publication of volumetric four-chamber meshes and volumetric bi-atrial meshes, respectively. Lastly, while [15] and [16] have studied the relationship between anatomical features, defined by their cardiac SSM, and in silico functional measurements of the heart, the impact of anatomical differences on functional measurements in real-world patients, which have the power to either validate these simulation findings or indicate modeling flaws, remains understudied.
@@ 1.4 Main contributions
Dataset construction: We use a large dataset comprising 271 high-resolution (submillimeter) CT images from a healthy, mixed-sex population, enhancing accuracy and antomical detail of the SSM. Moreover, fully automated segmentation and anatomical model generation workflows are implemented to build the SSM training dataset.
@@
Providing high-quality meshes for downstream applications: We provide a detailed biventricular SSM, together with a synthetic cohort of 100 ready-to-use biventricular meshes, including anatomical labels and fibers.
@@ 2.1.1 Raw input data
This study combines data from the Master@Heart study [28] with a proprietary dataset containing clinical data from female subjects, measured at the University Hospital of Leuven (UZ Leuven). The Master@Heart study is a multicenter (University Hospitals Leuven, University Hospital Antwerp, and Jessa Hospital Hasselt) prospective cohort trial, comparing cardiovascular health in lifelong endurance athletes, athletes who started training later in life, and healthy non-athletic controls, (ClinicalTrials.gov Identifier NCT03711539). In the remainder of this text, these subgroups will be referred to as early athletes, late athletes, and healthy controls, respectively. The Master@Heart dataset covers male subjects between the ages of 45 and 70. The following cardiovascular risk factors acted as exclusion criteria for this study: a medical history of cardiovascular disease, current or past smoking, use of antidiabetic drugs, statins, or antihypertensive drugs or a body mass index (BMI) exceeding . From this Master@Heart dataset, a subset, pertaining to subjects having their clinical measurements taken at UZ Leuven, was used. For consistency, female subjects of the same age group were selected from routine clinical care to complement this male-only dataset. Female subjects were excluded if they present with more than three of the following risk factors: smoking, arterial hypertension, diabetes, a family history of cardiovascular disease, and dyslipidemia. Given that this second dataset was obtained from clinical care rather than a dedicated clinical trial, the freedom in the selection process was more limited. However, the selection criteria are expected to select a population representative of healthy biventricular anatomy and function.
@@
| Early athletes | Male | 66 |
| Late athletes | Male | 57 |
@@
For both male and female subjects, ECG-triggered CT coronary angiography scans were acquired using a 128-slice dual-source CT scanner (Siemens Somatom Force, Siemens Healthineers, Forchheim, Germany), with images captured in diastole. The image resolutions are within the ranges for male subjects and for female subjects. The exact number of CT scans per demographic group is provided in Table 1. Ten-second, 12-lead ECG recordings were recorded for all subjects, except for three male subjects. The electrical axis was automatically determined from these ECGs using the Marquette 12SL ECG Analysis Program (GE Healthcare). Demographic data, including weight, height, sex, and age, and lifestyle data, indicating the subgroup (male early athletes, male late athletes, male healthy controls, or female subjects) were collected for all subjects, except for four male subjects. Secondary metrics, namely body mass index and body surface area (BSA) were calculated from weight and height. BSA was derived using the Du Bois formula [29].
@@ 2.2.1 Correspondence establishment and surface construction
For PCA, each anatomical geometry must be represented as a single feature vector, and comparability across different vectors, i.e. geometries, was established by means of corresponding point extractions from the original meshes. We tackled corresponding point extraction employing a new, lightweight method based on UVCs correspondence. To this end, a uniform sampling over the apico-basal , and rotational coordinate was applied to define a reference set of tuples . For both LV and RV, these tuples define a grid of corresponding points to be extracted from the endocardial () and epicardial () surfaces of the biventricular geometries. We sampled 80 values between for the apico-basal coordinate . For the rotational coordinate , we selected 160 values from the interval for the LV, and 80 values in the interval for the RV. A coarsening scheme was applied near the LV and RV apex, reducing the sampling density of to avoid having an increasingly dense point cloud in Cartesian space due to smaller ventricular circumference in the apical region, as illustrated in Figure 2. When , the sampling density of was decreased by factor 2, resulting in the selection of 80 values for the LV, and 40 for the RV. For , this sampling density was decreased by factor 4, selecting 40 and 20 values for the LV and RV, respectively. The sampling scheme resulted in point clouds consisting of points, comparable to the number of points included in meshes of similar studies [13, 14, 16].
@@ 2.2.2 Generalized Procrustes alignment of point clouds
The point clouds from all subjects were aligned using a generalized Procrustes method [43], an iterative technique that progressively moves the set of geometries, minimizing the Frobenius norm across all point clouds. In this iterative procedure, point clouds were translationally and rotationally aligned, while scaling was impeded, as size factors are intended to be captured in the SSM. The initial reference point cloud for the iterative procedure was randomly chosen from the set of point clouds, while it was defined as the average of all point clouds in each subsequent iteration. Rotation matrices for pairwise alignment of point clouds to the reference were calculated with the algorithm proposed by Schonemann [44], while translational movements were obtained from point cloud centering. The iterative procedure was automatically stopped at the stabilization of the average of all aligned point clouds, corresponding to a computed Frobenius norm between subsequent averages smaller than a prescribed tolerance , which was set in this work as .
@@ 2.2.3 SSM construction and evaluation
To investigate the reconstruction error of the SSM across different groups – namely, the female group, the male control group, male early athletes, and male late athletes – a leave-one-out cross-validation (LOOCV) experiment was performed over the full dataset. This experiment consisted of 271 iterations, equal to the total number of geometries. In each iteration, a single geometry was removed from the training set, and a new SSM was built from the remaining 270 geometries. This SSM was consequently used to reconstruct the left-out geometry, and the resulting reconstruction error was calculated. Comparing the reconstruction errors across the various demographic groups provides a notion of the suitability of the SSM to represent certain types of anatomies within this study population.
To investigate the number of geometries needed for effective SSM construction, we analyzed the reconstruction error across SSMs with varying training set sizes. Four training sets were constructed, containing either 56, 112, 168, or 224 geometries, randomly sampling an equal amount of geometries from each demographic group without replacement. A separate LOOCV experiment was carried out within each of these datasets.
Moreover, we tested the hypothesis that male and female cardiac anatomy differ significantly [45], by investigating the bias introduced by a sex-specific (biased) SSM. We quantified this bias as the reduced ability to represent geometries of the opposite sex compared to a mixed-sex (unbiased) SSM. To this end, we built two biased SSMs, one trained on 60 male subjects from the control group, and one trained on 60 healthy female subjects, in both cases randomly selected. We also trained an unbiased SSM on a random selection of 30 female and 30 male subjects, and we compared the main modes of variability of the three SSMs. For each dataset, we performed a LOOCV experiment to establish the baseline representation ability of each SSM. Moreover, we calculated the reconstruction error on geometries not present in the training set for each SSM. These were mixed-sex geometries for the unbiased SSM, and subjects from the opposite sex for both biased SSMs.
@@ 2.3.1 Synthetic surface mesh generation
We verified the underlying assumption of normality in A.2 for the first 16 PCs.
@@ 2.4.2 Statistical testing
To explore the relationship between biventricular anatomy and categorical demographic variables, mean differences between PC scores across demographic groups were computed. More specifically, the following comparisons were made: female subjects vs. male healthy controls, male healthy controls vs. male early athletes, male healthy controls vs. male late athletes, and male early athletes vs. male late athletes. The statistical significance of mean differences was evaluated using two-tailed t-tests.
@@ 3.2 Influence of demographics
Figure 8 illustrates the distribution of the reconstruction errors for the full SSM across the demographic groups included in this study. The mean reconstruction error for the female group was 0.54 , while those for the male control group, male early athletes, and male late athletes were 0.63 , 0.66 and 0.62 , respectively. The results show that the female group can be approximated with a slightly smaller reconstruction error. No substantial differences in reconstruction accuracy were obtained for the other groups, indicating that the full SSM is equally effective in representing biventricular geometries across the three demographic groups of the male population.
@@ 3.4 Impact of sex-related anatomical differences
Next, a quantitative analysis of the impact of sex-related anatomical differences on the SSM was conducted. The results of reconstructing male and female geometries on two sex-biased SSMs, and one unbiased SSM, are shown in Figure 10. In general, the reconstruction error was highest when reconstructing geometries with an opposite-sex SSM. For male geometries, this error was 1.48 , while for female geometries, it was 1.22 . Furthermore, similar performances were obtained when employing either an unbiased SSM or a biased SSM to reconstruct geometries of the same sex as the biased SSM used, i.e. 1.37 and 1.36 , respectively, for male subjects, and 1.16 and 1.13 , respectively, for female ones.

### v2

@@ Abstract
A cardiac digital twin is a virtual replica of a patient-specific heart, mimicking its anatomy and physiology. A crucial step of building a cardiac digital twin is anatomical twinning, where the computational mesh of the digital twin is tailored to the patient-specific cardiac anatomy. In a number of studies, the effect of anatomical variation on clinically relevant functional measurements like electrocardiograms (ECGs) is investigated, using computational simulations. While such a simulation environment provides researchers with a carefully controlled ground truth, the impact of anatomical differences on functional measurements in real-world patients remains understudied. In this study, we develop a biventricular statistical shape model and use it to quantify the effect of biventricular anatomy on ECG-derived and demographic features, providing novel insights for the development of digital twins of cardiac electrophysiology. To this end, a dataset comprising high-resolution cardiac CT scans from 271 healthy individuals, including athletes, is utilized. Furthermore, a novel, universal, ventricular coordinate-based method is developed to establish lightweight shape correspondence. The performance of the shape model is rigorously established, focusing on its dimensionality reduction capabilities and the training data requirements. The most important variability in healthy ventricles captured by the model is their size, followed by their elongation. These anatomical factors are found to significantly correlate with ECG-derived and demographic features. Additionally, a comprehensive synthetic cohort is made available, featuring ready-to-use biventricular meshes with fiber structures and anatomical region annotations. These meshes are well-suited for electrophysiological simulations.
@@ 1.1 Cardiac digital twins
Regardless of the specific aspect(s) of the cardiovascular system that a model targets, the anatomy of the individual needs to be taken into account. As such, defining a computational mesh representing this anatomy is a common step in the construction of all types of CDTs. This discrete, multi-label mesh is typically derived from cardiac computed tomography (CT) or magnetic resonance (MR) imaging of the subject. It delineates the computational domains for discretizing and numerically solving the PDEs governing the system. The process of creating a mesh that matches subject-specific anatomy is known as anatomical twinning [11].
On a broader scale, digital twins can be extended from individuals to populations, leading to the concept of virtual cohorts. These virtual cohorts enable in silico clinical trials that offer several advantages over traditional approaches, such as reducing the need for real patients to undergo experimental treatments, minimizing the risk of failed trials and associated financial costs, and optimizing trial design [12, 13]. Constructing an accurate anatomical representation of the target population is a crucial first step in the pipeline for developing functional, population-level CDTs.
Statistical shape modeling offers a powerful means to study and represent individual variation within a broader population framework [14, 15, 16, 17, 18, 19, 20]. Cardiac statistical shape models (SSMs), which capture population-level anatomical variations, closely relate to patient-level CDTs in the applied methodology, since both rely on anatomical twinning. Moreover, these models allow the creation of representative synthetic populations, enhancing virtual cohort studies and providing an anatomical building block for population-level CDTs.
@@ 1.2 Statistical shape modeling
(3) Statistical analysis: A cardiac SSM offers valuable, quantitative insights into heart shape variability, which can in turn be linked to other health-related data, such as electrophysiological measurements and demographic information, through statistical analyses [18, 19, 20]. As such, cardiac SSM are valuable tools for advancing our fundamental understanding of cardiac anatomy, notably in relation to other clinically relevant features, such as electrocardiograms (ECGs).
@@ 1.3 Related work
Previous work in cardiac SSM development provides meaningful insights into cardiac anatomy of various scopes (biventricular anatomy in [16, 17, 20], four-chamber cardiac anatomy in [18], and bi-atrial anatomy in [19]). However, some limitations can be identified. The spatial resolution of the input data is often limited compared to the clinical state-of-the-art [28, 16, 17, 19, 20]. Furthermore, some studies include only a limited and possibly insufficient number of subjects in their shape analysis, such as 19 shapes in [18] and 47 shapes in [19]. Such limitations are often related to the lack of automatic workflows for the generation of training geometries, resulting in time-consuming manual processes, especially in the segmentation step [17, 18, 19]. Moreover, some studies include subjects suffering from cardiovascular disease, meaning that the resulting SSM is likely not representative of a healthy population [17, 19, 20]. In terms of methodology, none of the studies leverage the recent advancements in the field of cardiac modeling, specifically designed for registration purposes, i.e. the implementation of universal cardiac coordinate systems [29, 30, 11, 31]. Additionally, although cardiac SSM studies usually include the release of synthetic cohorts, many of these cohorts only feature surface meshes [16, 17], and as such are not suitable for direct use in downstream applications, such as finite element simulations. Notable exceptions are [18] and [19], which include the publication of volumetric four-chamber meshes and volumetric bi-atrial meshes, respectively. Lastly, while [18] and [19] have studied the relationship between anatomical features, defined by their cardiac SSM, and in silico functional measurements of the heart, the impact of anatomical differences on functional measurements in real-world patients, which have the power to either validate these simulation findings or indicate modeling flaws, remains understudied.
@@ 1.4 Main contributions
Dataset construction: We use a large dataset comprising 271 high-resolution (submillimeter) CT images from a healthy, mixed-sex population, enhancing accuracy and anatomical detail of the SSM. Moreover, fully automated segmentation and anatomical model generation workflows are implemented to build the SSM training dataset. This automated anatomical twinning approach directly enables building patient-level, anatomical, biventricular CDTs.
@@
Providing high-quality meshes for downstream applications: We provide a detailed biventricular SSM, together with a synthetic cohort of 100 ready-to-use biventricular meshes, including anatomical labels and fibers, representative of healthy biventricular anatomy. The virtual anatomical cohort represents a building block towards a population-level CDT.
@@ 2.1.1 Raw input data
This study combines data from the Master@Heart study [32] with a proprietary dataset containing clinical data from female subjects, measured at the University Hospital of Leuven (UZ Leuven). All of these subjects were screened for absence of coronary artery disease. The Master@Heart study is a multicenter (University Hospitals Leuven, University Hospital Antwerp, and Jessa Hospital Hasselt) prospective cohort trial, comparing cardiovascular health in lifelong endurance athletes, athletes who started training later in life, and healthy non-athletic controls, (ClinicalTrials.gov Identifier NCT03711539). In the remainder of this text, these subgroups will be referred to as early-onset athletes, late-onset athletes, and healthy controls, respectively. The Master@Heart dataset covers male subjects between the ages of 45 and 70. The following cardiovascular risk factors acted as exclusion criteria for this study: a medical history of cardiovascular disease, current or past smoking, use of antidiabetic drugs, statins, or antihypertensive drugs or a body mass index (BMI) exceeding 27.2 . From this Master@Heart dataset, a subset, pertaining to subjects having their clinical measurements taken at UZ Leuven, was used. For consistency, female subjects of the same age group were selected from routine clinical care to complement this male-only dataset. Female subjects were excluded if they present with more than three of the following risk factors: smoking, arterial hypertension, diabetes, a family history of cardiovascular disease, and dyslipidemia. Given that this second dataset was obtained from clinical care rather than a dedicated clinical trial, the freedom in the selection process was more limited. However, the selection criteria are expected to select a population representative of healthy biventricular anatomy and function.
@@
| Early-onset athletes | Male | 66 |
| Late-onset athletes | Male | 57 |
@@
For both male and female subjects, ECG-triggered CT coronary angiography scans were acquired using a 128-slice dual-source CT scanner (Siemens Somatom Force, Siemens Healthineers, Forchheim, Germany), with images captured in diastole. The image resolutions are within the ranges for male subjects and for female subjects. The exact number of CT scans per demographic group is provided in Table 1. Ten-second, 12-lead ECG recordings were recorded for all subjects, except for three male subjects. The electrical axis was automatically determined from these ECGs using the Marquette 12SL ECG Analysis Program (GE Healthcare). Demographic data, including weight, height, sex, and age, and lifestyle data, indicating the subgroup (male early-onset athletes, male late-onset athletes, male healthy controls, or female subjects) were collected for all subjects, except for four male subjects. Secondary metrics, namely body mass index and body surface area (BSA) were calculated from weight and height. BSA was derived using the Du Bois formula [33].
@@ 2.2.1 Correspondence establishment and surface construction
For PCA, each anatomical geometry must be represented as a single feature vector, and comparability across different vectors, i.e. geometries, was established by means of corresponding point extractions from the original meshes. We tackled corresponding point extraction employing a new, lightweight method based on UVCs correspondence. To this end, a uniform sampling over the apico-basal , and rotational coordinate was applied to define a reference set of tuples . For both LV and RV, these tuples define a grid of corresponding points to be extracted from the endocardial () and epicardial () surfaces of the biventricular geometries. We sampled 80 values between for the apico-basal coordinate . For the rotational coordinate , we selected 160 values from the interval for the LV, and 80 values in the interval for the RV. A coarsening scheme was applied near the LV and RV apex, reducing the sampling density of to avoid having an increasingly dense point cloud in Cartesian space due to smaller ventricular circumference in the apical region, as illustrated in Figure 2. When , the sampling density of was decreased by factor 2, resulting in the selection of 80 values for the LV, and 40 for the RV. For , this sampling density was decreased by factor 4, selecting 40 and 20 values for the LV and RV, respectively. The sampling scheme resulted in point clouds consisting of points, comparable to the number of points included in meshes of similar studies [16, 19]. The point extraction process is illustrated in Figure 3. For the rationale behind selecting this sampling scheme, we refer to A.
@@ 2.2.2 Generalized Procrustes alignment of point clouds
The point clouds from all subjects were aligned using a generalized Procrustes method [47], an iterative technique that progressively moves the set of geometries, minimizing the Frobenius norm across all point clouds. In this iterative procedure, point clouds were translationally and rotationally aligned, while scaling was impeded, as size factors are intended to be captured in the SSM. In a parallel analysis, we performed a Procrustes alignment where scaling was allowed, to study heart shape effects independently of heart size, as explained in C. The initial reference point cloud for the iterative procedure was randomly chosen from the set of point clouds, while it was defined as the average of all point clouds in each subsequent iteration. Rotation matrices for pairwise alignment of point clouds to the reference were calculated with the algorithm proposed by Schonemann [48], while translational movements were obtained from point cloud centering. The iterative procedure was automatically stopped at the stabilization of the average of all aligned point clouds, corresponding to a computed Frobenius norm between subsequent averages smaller than a prescribed tolerance , which was set in this work as .
@@ 2.2.3 SSM construction and evaluation
To investigate the reconstruction error of the SSM across different groups – namely, the female group, the male control group, male early-onset athletes, and male late-onset athletes – a leave-one-out cross-validation (LOOCV) experiment was performed over the complete dataset on the one hand, and a stratified dataset, containing 56 subjects from each of the four demographic subgroups, on the other hand. The two experiments consisted of 271 and 224 iterations, respectively, equal to the total number of geometries in the dataset of the experiment. In each iteration, a single geometry was removed from the training set, and a new SSM was built from the remaining geometries. This SSM was consequently used to reconstruct the left-out geometry, and the resulting reconstruction error was calculated. Comparing the reconstruction errors across the various demographic groups provides a notion of the suitability of the SSM to represent certain types of anatomies within this study population.
To investigate the number of geometries needed for effective SSM construction, we analyzed the reconstruction error across SSMs with varying training set sizes. Four training sets were constructed, containing either 56, 112, 168, or 224 geometries, randomly sampling an equal amount of geometries from each demographic group without replacement. A fifth dataset contained all 271 available geometries. A separate LOOCV experiment was carried out within each of these five datasets.
Furthermore, we tested the hypothesis that male and female cardiac anatomy differ significantly [49], by investigating the bias introduced by a sex-specific (biased) SSM. We quantified this bias as the reduced ability to represent geometries of the opposite sex compared to a mixed-sex (unbiased) SSM. To this end, we built two biased SSMs, one trained on 60 male subjects from the control group, and one trained on 60 healthy female subjects, in both cases randomly selected. We also trained an unbiased SSM on a random selection of 30 female and 30 male subjects, and we compared the main modes of variability of the three SSMs. For each dataset, we performed a LOOCV experiment to establish the baseline representation ability of each SSM. Moreover, we calculated the reconstruction error on geometries not present in the training set for each SSM. These were mixed-sex geometries for the unbiased SSM, and subjects from the opposite sex for both biased SSMs.
@@ 2.3.1 Synthetic surface mesh generation
We verified the underlying assumption of normality in E for the first 16 PCs.
@@ 2.4.2 Statistical testing
To explore the relationship between biventricular anatomy and categorical demographic variables, mean differences between PC scores across demographic groups were computed. More specifically, the following comparisons were made: female subjects vs. male healthy controls, male healthy controls vs. male early-onset athletes, male healthy controls vs. male late-onset athletes, and male early-onset athletes vs. male late-onset athletes. The statistical significance of mean differences was evaluated using two-tailed t-tests.
@@ 3.2 Influence of demographics
Figure 9 illustrates the distribution of the reconstruction errors for the full SSM across the demographic groups included in this study. The mean reconstruction error for the female group was 0.54 , while those for the male control group, male early-onset athletes, and male late-onset athletes were 0.63 , 0.66 and 0.62 , respectively. The results show that the female group can be approximated with a slightly smaller reconstruction error. No substantial differences in reconstruction accuracy were obtained for the other groups, indicating that the full SSM is equally effective in representing biventricular geometries across the three demographic groups of the male population. An additional analysis in B shows that this effect is not caused by the slight demographic imbalance in the dataset, and C shows that the smaller reconstruction errors associated with female harts can largely, but not entirely be explained by heart size factors.
@@ 3.4 Impact of sex-related anatomical differences
Next, a quantitative analysis of the impact of sex-related anatomical differences on the SSM was conducted. The results of reconstructing male and female geometries on two sex-biased SSMs, and one unbiased SSM, are shown in Figure 11. For the reconstruction of male geometries, the reconstruction error was highest, namely 1.48 , when using an opposite-sex SSM. The reconstruction error when using a same- or mixed-sex SSM was lower, respectively 1.36 and 1.37 . The same holds for female geometries, where an opposite-, same- and mixed-sex SSM leads to reconstruction errors equal to 1.22, 1.13, and 1.16 , respectively. For an additional analysis of the influence of sex-related anatomical differences on SSM representativeness, independent of heart size, we refer to C.